How do you find arcsin(sin3) ??
Where 3 is in radians .

Grade:11

1 Answers

bharat bajaj IIT Delhi
10 years ago
You know 3 comes between pi/2 and pi.
Our answer should be in between -pi/2,pi/2.
sin3 = sin ( pi - (pi-3) ) = sin(pi-3)
Hence, arcsin(sin3) = arcsin(sin(pi-3)) = pi-3
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty